# Generalized Sato-Tate and quadratic residues

### v1

@@ 1. Introduction
The present note is a continuation of [KTVZ] and treats some questions on the distribution of consecutive quadratic residues, left there without answer.
One the main results of [KTVZ] is rather exact information on the distribution of consecutive quadratic residues modulo very large primes. Those results are subject to strong statistical independence conditions which seem natural, but completely out of reach.
Our main observation in the present note is that those conditions follow from some particular cases of the Generalized Sato-Tate conjecture (GST) and that some already known cases of GST lead to some interesting bounds (both upper and lower) concerning corresponding arithmetic functions.
The rest of the note is organized as follows.
In Section 2 we recall the formulation of GST, in Section 3 we cite some facts from [KTVZ] concerning consecutive quadratic residues and their connections with algebraic curves and abelian varieties over finite fields.
Section 4 contains some asymptotic results on the distribution of 4 and 5 consecutive quadratic residues under validity of GST for certain products of elliptic curves of dimension 5 or less.
In Section 5 we prove unconditionally some bounds similar to, but weaker than those of Section 4 using known cases of GST for products of 2 elliptic curves.
@@ 2. The Generalized Sato-Tate conjecture
b). $\mathrm{End}_{\mathbb{C}}(E)=\mathbb{Z}$, that is, $E$ has no complex multiplication, and $\mathrm{ST}_{E}=\mathrm{SU}_{2}(\mathbb{C}).$ Note that in any case $\mathrm{End}_{\mathbb{Q}}(E)=\mathbb{Z}.$
@@
However, the next case of interest for us has $g=17$ and thus is rather far from the current proof methods.
Nevertheless, in the rest of our note we suppose, except in Section 5 that GST(A) is true for all abelian varieties mentioned below.
@@ 3. Consecutive quadratic residues and algebraic curves
Below we will concentrate only on the behaviour of $n_{p}(R^{t})$ and denote it by $n_{p}(t)$. For other cases and some detail see our preceding paper [KTVZ].
@@
of genus $g_{T}=g(C_{T})=\lfloor(s-1)/2\rfloor$ for $s=|T|,$ and $p+1-N_{T}$, is just the trace of Frobenius for $C_{T}$.
@@
$\bf{t=2}.$ Here $g_{2}=0,$ there are 3 possible $T,$ all with rational $C_{T},\,N_{T}=0$ and $c_{p}=-\frac{1}{4}$ or
$c_{p}=-\frac{3}{4}$ depending on p(mod 4).p(mod\;4).
@@
which leads to Jacobsthal’s formula with $c_{p},|c_{p}|\leq\frac{3}{2}$ depending on p(mod 8).p(mod\;8).
@@
$\bf{t=6}.$ If $t=6\;g_{6}=49$, then some genus 2 curves $C_{T}$ for $|T|=5$ have simple Jacobians, which leads to an expression of $n_{p}(6)$ in terms of Frobenius traces for numerous ($\geq 35$) elliptic curves and few ($\leq 7$) curves of genus 2.
@@ 4.2. Case $t=4$
$\bar{E}_{1}$ being the non-trivial twist of $E_{1}$ over $\mathbb{Q}(i)$. Assuming GST for $A(4),$ we get
$\mathrm{ST(A(4))}\subseteq\Delta_{1}\times N\times\mathrm{SU_{2}(\mathbb{C})}\subset\mathrm{\bf{T}}\times\mathrm{\bf{T}}\times\mathrm{SU_{4}(\mathbb{C})}\times\mathrm{SU_{2}(\mathbb{C})}\subset\mathrm{SU_{10}(\mathbb{C})},$
where $\bf{T}=\mathrm{U_{1}(\mathbb{C})}\subset\mathrm{SU_{2}(\mathbb{C})},$ $\Delta_{1}$ being the image of the diagonal embedding $\bf{T}\hookrightarrow\mathrm{SU_{4}(\mathbb{C})}$, whereas $N$ is the normalizer of $\mathrm{SU_{2}(\mathbb{C})}\hookrightarrow\mathrm{SU_{4}(\mathbb{C})}$. In (5) and (6) the summands $0,0,a_{4}(p)/16$ and $a_{0}(p)/8,a_{1}(p)/8,a_{4}(p)/16,$ respectively, correspond to the measures
@@
$(ii)$ Assume GST for the product $A=E_{0}\times E_{1}\times E_{4}.$ Then the quantity $\delta_{p}(4)$ for $p$ running the set $\mathcal{P}_{1}:=\{p=4m+1\;\;is\;\ prime\}$ is equidistributed on [10,10][-10,10] with respect to the measure
@@
It follows directly from formulas (1),(8) and (6).
Therefore
@@
$h_{11}:=\frac{\sqrt{16-s^{2}}*\sqrt{4-u^{2}}}{32\pi^{2}}\geq 0,\;h_{12}=\frac{1}{16\pi^{3}\sqrt{16-v^{2}}}*\sqrt{16-s^{2}}*\sqrt{4-u^{2}}\geq 0.$
@@
Remark. We have used that $E_{0},E_{1},E_{4}$ are not isogenous.
Indeed, $E_{0},E_{1}$ and $E_{4}$ are pairly non-isogenous (over $\bar{\mathbb{Q}}$). It is clear for $(E_{0},E_{1})$ and $(E_{0},E_{2})$ since $E_{0}$ has CM, while $E_{1},E_{2}$ do not. For the pair $(E_{1},E_{2})$ it is sufficient to look at their reductions modulo 7 to see that $E^{\prime\prime}=E_{2}\mod 7$ is supersingular, $a_{2}(7)=0,$ whereas $E^{\prime}=E_{1}\mod 7$ is not, $a_{1}(7)=-4.$ Thus $E^{\prime}$ and $E^{\prime\prime}$ are not isogenous over $\bar{\mathbb{F}}_{7}$ and hence $E_{1},E_{2}$ are not isogenous over $\bar{\mathbb{Q}}$. ∎
Note that this Theorem 4.1 implies that the sequences $\{a_{0}(p)\},\{a_{1}(p)\}$ and $\{a_{4}(p)\}$ are statistically independent.
@@ 4.3. Case $t=5$
the curve $C$ admits a non-hyper-elliptic involution
@@
Admitting GST for $A(5),$ one gets a result parallel to Theorem 4.1 where the formulas are rather lengthy and include numerous convolutions; in fact it is sufficient to admit GST for
$B(5):=E_{0}\times E_{1}\times E_{4}\times E_{12}\times E_{15},\;\dim B(5)=5.$
@@
In general we get, using the splitting rule of $p$ in $\mathbb{Q}(\sqrt{-1})$ and $\mathbb{Q}(\sqrt{2})$:
@@ 5. Unconditional results
$p_{1}=p_{1}(\varepsilon)\equiv 1\mod 4,\;p^{\prime}_{1}=p_{1}^{\prime}(\varepsilon)\equiv 1\mod 4,$
@@
and since $|\frac{a_{4}(p_{1})}{\sqrt{p_{1}}}|\leq 2$, an immediate calculation implies the conclusion.∎
@@
The same GST case for $A=E\times F$ can also be applied to get unconditional results for $t=5$ in the spirit of Proposition 5.2 which are rather weak, yet non-trivial.
@@
$p_{1}=p_{1}(\varepsilon)\equiv 1(8)=p_{3}^{\prime}(\varepsilon)\equiv 1(8),$

### v2

@@ 1. Introduction
The present note is a continuation of [KTVZ] and deals with several questions concerning the distribution of consecutive quadratic residues that were left open there.
One of the main results of [KTVZ] provides rather precise information on the distribution of consecutive quadratic residues modulo very large primes. These results are based on strong statistical independence assumptions that appear natural but are, at present, completely out of reach.
The main observation in this note is that these assumptions follow from certain particular cases of the Generalized Sato–Tate conjecture (GST). Moreover, some already established cases of the GST yield interesting upper and lower bounds for the corresponding arithmetic functions.
Section 2 recalls the formulation of the GST. Section 3 summarizes several results from [KTVZ] concerning consecutive quadratic residues and their connections with algebraic curves and abelian varieties over finite fields. Section 4 presents asymptotic results on the distribution of four and five consecutive quadratic residues, under the assumption of the GST for certain products of elliptic curves of dimension at most five. Finally, Section 5 establishes unconditional bounds, similar in form but weaker than those of Section 4, by using known cases of the GST for products of two elliptic curves.
@@ 2. The Generalized Sato-Tate conjecture
b). $\mathrm{End}_{\mathbb{C}}(E)=\mathbb{Z}$, that is, $E$ has no complex multiplication, and $\mathrm{ST}_{E}=\mathrm{SU}_{2}(\mathbb{C}).$
Note that in any case $\mathrm{End}_{\mathbb{Q}}(E)=\mathbb{Z}.$
@@
However, the next case of interest for us has $g=17$ and thus is far from the current proof methods.
Nevertheless, in the rest of our note we suppose, except in Section 5 that GST(A) is true for all abelian varieties $A$ used below.
@@ 3. Consecutive quadratic residues and algebraic curves
Below we will concentrate only on the behaviour of $n_{p}(R^{t})$ and denote it by $n_{p}(t)$. For other cases and some detail see our paper [KTVZ].
@@
of genus $g_{T}=g(C_{T})=\lfloor(s-1)/2\rfloor$ for $s=|T|,$ then $p+1-N_{T}$ is just the trace of Frobenius for $C_{T}$.
@@
$\bf{t=2}.$ Here $g_{2}=0,$ there are 3 possible $T,$ all with rational $C_{T},\,N_{T}=0$ and $c_{p}=-\frac{1}{4}$ or $c_{p}=-\frac{3}{4}$ depending on $p\mod 4.$
@@
which leads to Jacobsthal’s formula with $c_{p},|c_{p}|\leq\frac{3}{2}$ depending on $p\mod 8.$
@@
$\bf{t=6}.$ If $t=6,\;g_{6}=49$, then some genus 2 curves $C_{T}$ for $|T|=5$ have simple Jacobians, which leads to an expression of $n_{p}(6)$ in terms of Frobenius traces for numerous ($\geq 35$) elliptic curves and few ($\leq 7$) curves of genus 2.
@@ 4.2. Case $t=4$
$\bar{E}_{1}$ being the non-trivial twist of $E_{1}$ over $\mathbb{Q}(i)$.
Assuming GST for $B(4):=E_{0}\times E_{1}\times E_{4}$, we get
$\mathrm{ST(B(4))}\subseteq\Delta_{1}\times N\times\mathrm{SU_{2}(\mathbb{C})}\subset\mathrm{\bf{T}}\times\mathrm{SU_{2}(\mathbb{C})}\times\mathrm{SU_{2}(\mathbb{C})}\subset\mathrm{SU_{6}(\mathbb{C})},$
where $\bf{T}=\mathrm{U_{1}(\mathbb{C})}\subset\mathrm{SU_{2}(\mathbb{C})},$ $\Delta_{1}$ being the image of the diagonal embedding $\bf{T}\hookrightarrow\mathrm{SU_{2}(\mathbb{C})}$, whereas $N$ is the normalizer of $\bf{T}\hookrightarrow\mathrm{SU_{2}(\mathbb{C})}$. In (5) and (6) the summands $0,0,a_{4}(p)/16$ and $a_{0}(p)/8,a_{1}(p)/8,a_{4}(p)/16,$ respectively, correspond to the measures
@@
$(ii)$ Assume GST for $B(4)=E_{0}\times E_{1}\times E_{4}.$ Then the quantity $\delta_{p}(4)$ for $p$ running the set $\mathcal{P}_{1}:=\{p=4m+1\;\;is\;\ prime\}$ is equidistributed on [10,10][-10,10] with respect to the measure
@@
The proof follows directly from formulas (1),(8) and (6).
Therefore,
@@
$h_{11}:=\frac{\sqrt{16-s^{2}}*\sqrt{4-u^{2}}}{32\pi^{2}}\geq 0,\;h_{12}:=\frac{1}{16\pi^{3}\sqrt{16-v^{2}}}*\sqrt{16-s^{2}}*\sqrt{4-u^{2}}\geq 0.$
@@
Remark. We have used that $E_{0},E_{1},E_{4}$ are not pairwise isogenous.
Clearly, $E_{0}$ has CM, and $E_{1},\bar{E}_{1},E_{4}$ have not, and thus $E_{0}$ is not isogenous (over $\bar{\mathbb{Q}}$) to $E_{1},\bar{E}_{1}$ or $E_{4};$ also, $E_{1},\bar{E}_{1}$ are not isogenous to $E_{4}$ , since $tr(E^{\prime}_{1}),tr(\bar{E}^{\prime}_{1})=\pm 2$ and $trE_{4}^{\prime}=-1$ for $E^{\prime}_{k}:=E_{k}\mod 5,$ and the traces are not equal up to sign. ∎
Note that GST($B(4)$) implies that the sequences $\{a_{0}(p)\},\{a_{1}(p)\}$ and $\{a_{4}(p)\}$ are statistically independent.
@@ 4.3. Case $t=5$
note that the curve $C$ admits a non-hyper-elliptic involution
@@
Indeed, $E_{0},E_{1}$ and $E_{4}$ are pairly non-isogenous (over $\bar{\mathbb{Q}}$). For the pair $(E_{1},E_{12})$ it is sufficient to look at their reductions modulo 7 to see that $E^{\prime\prime}_{1}=E_{1}\mod 7$ is supersingular, $a_{1}(7)=0,$ whereas $E^{\prime\prime}_{12}=E_{12}\mod 7$ is not, $a_{12}(7)=-1.$ Thus $E_{1}^{\prime\prime}$ and $E^{\prime\prime}_{12}$ are not isogenous over $\bar{\mathbb{F}}_{7}$ and hence $E_{1},E_{12}$ are not isogenous over $\bar{\mathbb{Q}}$. The same for $(E_{1},E_{16})$ since $a_{16}(7)=5.$ For $(E_{4},E_{12}):a_{4}(11)=5,a_{12}(11)=-3;$ for $(E_{4},E_{16}):a_{4}(7)=1,a_{16}(7)=5;$ finally, $a_{12}(7)=-1,a_{16}(7)=5$ gives the same for $(E_{12},E_{16})$.
Admitting GST for $A(5),$ one gets a result parallel to Theorem 4.1 where the formulas are more lengthy and include more convolutions; in fact it is sufficient to admit GST for
$B(5):=E_{0}\times E_{1}\times E_{4}\times E_{12}\times E_{16},\;\dim B(5)=5.$
@@
In general we get, using how $p$ splits in $\mathbb{Q}(\sqrt{-1})$ and $\mathbb{Q}(\sqrt{2})$:
@@ 5. Unconditional results
$p_{1}=p_{1}(\varepsilon)\equiv 1\mod 4,\;p_{1}=p_{1}^{\prime}(\varepsilon)\equiv 1\mod 4,$
@@
and since $|{a_{4}(p_{1})}|\leq 2\sqrt{p_{1}}$, an immediate calculation implies the conclusion.∎
@@
[…]
@@
$p_{1}=p_{1}(\varepsilon)\equiv 1\mod 8,\;p_{1}=p_{1}^{\prime}(\varepsilon)\equiv 1\mod 8,$